OpenStudy (anonymous):

need help, inverse functions Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21. −45 0 2 63

OpenStudy (anonymous):

can you help?

OpenStudy (radar):

f(x)=y y = 6(x+2)-3 Swap the x and y x = 6(y+2)-3 x= 6y + 12 -3 x= 6y + 9 Isolate the y term 6y =x-9 divide both sides by 6 y =(x-9)/6 Now substitute 21 for x as stated in the problem. y =(21-9)/6 y = 12/6 = 2 y=2

OpenStudy (radar):

or f(x) inverse = 2 when x=21

OpenStudy (anonymous):

Ok, i get it now. thanks!

OpenStudy (radar):

You're welcome and good luck with your studies.